### v2

# DMRetriever: A Family of Models for Improved Text Retrieval in Disaster Management

Abstract
[…]
DMRetriever: A Family of Models for Improved Text Retrieval in Disaster Management
Kai Yin1   Xiangjue Dong1   Chengkai Liu1   Allen Lin1 Lingfeng Shi1Ali Mostafavi1James Caverlee1 1Texas A&M University {kai_yin,xj.dong,liuchengkai,al001,lingfengs111,mostafavi,caverlee}@tamu.edu
1 Introduction
Disasters, both natural and man-made, inevitably disrupt communities, endanger lives, and cause severe economic losses (fan2021disaster; liu2025artificial). During such events, emergency responders, policymakers, and affected populations rely on information retrieval (IR) systems for decision-making, situational awareness, and coordination of relief efforts (abbas2025exploring). Timely and reliable information access is thus essential for effective disaster management (priya2020taqe).
Figure 1: Average performance of DMRetriever. Breakdown for each search intent is in Table 3.4.3
Information needs during disaster events are highly diverse, spanning fact-checking critical information, monitoring social media, and consulting technical or policy documents (See §3.1 for full lists of search intents) (yin2025disastir). However, a recent analysis of 30 popular general-domain IR models finds that none consistently achieve SOTA performance across all disaster management-specific search intents (yin2025disastir). This limits their real-world applicability: some queries yield valuable insights, while others return irrelevant results, delaying decision-making and risking missed evidence that could compromise response efforts.
Furthermore, real-world disaster management demands scalable retrieval solutions, from lightweight models for resource-constrained settings (e.g., on edge devices for rapid response) to larger models for performance-critical decision-making (e.g., in central command system). Hence, there is a great opportunity to develop domain-specialized IR models of various scales that achieve both high performance and computational efficiency for disaster management.
Figure 2: Overview of data refinement pipeline and progressive fine-tuning for DMRetriever. MTP: Massive Text Pairs, MTT: Massive Text Triplets. Both MTP and MTT in sub-figure (B) include DM (Disaster Management) and GD (General Domain) variants, with constructions of GD variants omitted here and detailed in §3.3.1 and §3.4.3. DM passages include those extracted from PDFs and those generated by LLMs. DMRetriever-Init: backbone after bidirectional attention adaptation for decoder-only; original backbone for encoder-only. DMRetriever-PT: model after pre-training. DMRetriever-It2: model after iteration 2. Algorithm 1 details the full workflow.
To this end, we introduce DMRetriever, a family of dense retrieval models at six scales (33M–7.6B parameters) to improve retrieval performance in disaster management (Figure 1).
The success of domain-specific IR models in fields such as biomedicine (xu2024bmretriever), law (chalkidis2020legal), and finance (sarmah2024hybridrag) shows the effectiveness of adding deep domain knowledge to improve performance. However, injecting disaster management knowledge into DMRetriever to improve performance poses two challenges: (1) the lack of high-quality, knowledge-rich data that covers diverse search intents during disasters, and (2) the absence of an efficient training method that can transfer such knowledge across models of different sizes.
To tackle the first challenge, we propose a data refinement pipeline (A in Figure 2) to generate the first, knowledge-rich data for disaster management. We first prompt a large language model (LLM) [^1] We use GPT-4o-mini as LLM in this work. with domain-specific passages to generate large-scale unlabeled query–passage pairs [^2] Following xiao2024c; wang2022text, we treat data without mined negatives as unlabeled and data with mined hard negatives as labeled. (A-1 in Figure 2, §3.3.1). Given limitations of existing IR models in disaster management, we refine these noisy corpora with mutual-agreement filtering, which removes mislabeled pairs through model consensus (A-2 in Figure 2, §3.4.2). Recognizing the varied learning capacities of different-sized DMRetriever models, we further propose difficulty-aware hard negative mining to generate negatives at varying difficulty levels (A-3 in Figure 2, §3.4.3), finally yielding a high-quality labeled dataset.
For the second challenge, we propose a Progressive instruction fine-tuning (B in Figure 2) that enables DMRetriever models of different sizes to effectively learn domain knowledge. For large variants initialized from decoder-only backbones (Table 1), we adapt their causal attention into bidirectional attention, allowing access to information from future tokens (B-1 in Figure 2, §3.2). Next, to better initialize DMRetriever and inject domain knowledge, we perform large-scale contrastive pre-training on unlabeled query–passage pairs (B-2 in Figure 2, §3.3). Finally, to adapt DMRetriever to diverse search intents, we perform instruction fine-tuning on labeled data2, where for small variants that have limited learning capacity, the difficulty of negative samples is progressively increased during fine-tuning (B-3 in Figure 2, §3.4.4).
[…]
The contributions of this work are as follows:
  • (1)
    We introduce DMRetriever, a family of six dense retrieval models (33M to 7.6B) that achieve new SOTA performance on each search intent across all scales, with superior parameter efficiency over existing baselines.
    Model #Parameters Backbone Backbone Type Hidden Size #Layers
    DMRetriever-33M 33M MiniLM encoder-only 384 12
    DMRetriever-109M 109M BERT-base-uncased encoder-only 768 12
    DMRetriever-335M 335M BERT-large-uncased-WWM encoder-only 1024 24
    DMRetriever-596M 596M Qwen3-0.6B decoder-only 1024 28
    DMRetriever-4B 4B Qwen3-4B decoder-only 2560 36
    DMRetriever-7.6B 7.6B Qwen3-8B decoder-only 4096 36
    Table 1: Model configurations of DMRetriever family across six model sizes. DMRetriever-7.6B is initialized from Qwen3-8B after removing its language modeling head, leaving 7.568B parameters (rounded to 7.6B).
  • (2)
    We propose a data refinement pipeline yielding the first large-scale training dataset for disaster management. It is generalizable to other domains with limited annotations.
  • (3)
    We introduce progressive instruction fine-tuning which enables models of varied sizes to effectively absorb domain knowledge and enhances their ability across different search intents.
2 Related Work
General-domain retrieval has been long dominated by bidirectional encoders such as BERT (devlin2019bert) and T5 (raffel2020exploring), with models like Sentence-BERT (reimers2019sentence) and SimCSE (gao2021simcse) improving performance through contrastive pre-training and supervised fine-tuning on MS MARCO (nguyen2016ms). More recently, decoder-only models, from GPT-based embeddings (brown2020language) to models such as Qwen3-Embedding (qwen3embedding), have leveraged instruction tuning and large-scale synthetic data to achieve SOTA results.
However, these general-domain models face clear limitations in disaster management, with no approach achieving SOTA performance across diverse search intents (yin2025disastir) due to distribution shift issue (thakur2021beir). Prior work on disaster information access has mainly targeted on social media, emphasizing clustering of posts, event or situational classification, and named entity recognition (imran2015processing; alam2021crisisbench; yin2024crisissense). While valuable, these efforts remain narrow in scope and have yet to address the need for robust retrieval systems tailored to disaster management-related diver search intents.
3 DMRetriever
DMRetriever is a family of dense retrieval models at six scales (33M–7.6B) initialized from encoder- and decoder-only backbones (Table 1). This scalability supports real-world disaster management applications ranging from resource-constrained settings with smaller models to performance-critical scenarios with larger ones.
3.1 Preliminaries
In DMRetriever, queries and passages are encoded into dense vectors using a shared embedding model $\mathbf{E}(\cdot)$. Given a query $q$ and a passage $p$, we prepend a search intent-specific instruction $I_{q}$ to the query to improve search intent awareness (wei2021finetuned; ouyang2022training; lee2024nv).
Following DisastIR (yin2025disastir), we consider six disaster management-specific search intents: question-answer (QA), Twitter (TW), Fact Checking (FC), Natural Language Inference (NLI), and Semantic Textual Similarity (STS). Table 5 details corresponding instruction formats.
The embeddings are then computed as $\mathbf{e}_{q}=\mathbf{E}(I_{q}\oplus q)$ and $\mathbf{e}_{p}=\mathbf{E}(p)$, where $\oplus$ denotes sequence concatenation. We apply mean pooling over the final hidden states to obtain fixed-length representations. The relevance score $\text{sim}(q,p)$ is computed as the cosine similarity between embeddings, scaled by a temperature parameter $\tau$:
$\text{sim}(q,p)=\cos(\mathbf{e}_{q},\mathbf{e}_{p})/\tau.$ (1)
To effectively adapt DMRetriever to disaster management domain, we propose a three-stage training framework using high-quality training data produced from our advanced data refinement pipeline (Figure 2 and Algorithm 1).
3.2 Enabling Bidirectional Attention in Decoder-only Backbones
3.2.1 Bidirectional attention adaptation
For DMRetriever initialized from decoder-only backbones, we replace the causal attention mask with an all-ones matrix (B-1 in Figure 2) to enable the model to capture information from future tokens, inspired by the recent success of lee2024nv; behnamghader2024llm2vec; muennighoff2024generative. This replacement allows each token to attend to every other token in the sequence, converting it into a bidirectional model.
To make model aware of its bidirectional attention, we train model by using masked language model (MLM) training objective (devlin2019bert). Given an input sequence, we randomly mask some of tokens and then train model to predict masked tokens based on past and future tokens.
3.2.2 MLM training data construction
We perform MLM training using a mix of general-domain and disaster management corpora. For the general domain, we use the English Wikipedia from merity2016pointer. For the disaster management domain, we use passages from DM-MTP (constructed in §3.3.1).
3.3 Unsupervised contrastive pre-training
3.3.1 Pre-training dataset construction
To robustly initialize DMRetriever for retrieval tasks and inject domain knowledge, we construct MTP (Massive Text Pairs; see Algorithm 2), which consists of two large-scale subsets: DM-MTP (disaster management domain) and GD-MTP (general-domain). Its composition is shown in Table 10.
In the absence of paired datasets for disaster management, DM-MTP represents the first large-scale collection of query–passage pairs in this domain. Inspired by wang2023improving and yin2025disastir, we propose a three-stage pipeline to generate it (A-1 in Figure 2, lines 1–7 in Algorithm 2). In the first stage, we crawl the web with queries formed by combining disaster event type names and location names, collect domain-specific PDFs, and process them into chunks, taking these as disaster-domain passages. Secondly, for each passage, a LLM generates information-need statement taking passage content and its designated search intent definition as input. Finally, the information-need statement is paired with the passage to prompt the LLM to generate a user query and a directly relevant passage, which serves as the positive passage for this query in DM-MTP. Appendix A details the whole DM-MTP construction process.
To enrich DMRetriever with diverse linguistic and semantic relevance patterns, GD-MTP is built by aggregating 15 publicly available datasets (Table 10). Query–positive passage pairs are created from unlabeled data using search intent-specific heuristics (Lines 8–12 in Algorithm 2). For example, in datasets associated with QA intents, user questions are treated as queries and their answers as positive passages. Heuristics for other search intents are described in Appendix B.
For negative sampling in MTP, instead of explicitly mining hard negatives, we adopt the in-batch negative strategy (karpukhin2020dense), where passages from other pairs within the same mini-batch serve as negative examples.
3.3.2 Contrastive Pre-training
DMRetriever is pre-trained on MTP using contrastive learning to separate relevant query–passage pairs from irrelevant ones. For each mini-batch $\mathcal{B}$, we adopt InfoNCE loss (chen2020simple) as pre-training objective, encouraging similarity between $q_{i}$ and its positive passage $p_{i}^{+}$ to be higher than that with any other passage $p_{j}\in\mathcal{B},\,j\neq i$ within the same batch (B-2 in Figure 2):
$\mathcal{L}_{\text{cpt}}=-\log\frac{e^{\mathrm{sim}(q_{i},p_{i}^{+})}}{\sum\limits_{p_{j}\in\mathcal{B}}e^{\mathrm{sim}(q_{i},p_{j})}}.$ (2)
3.4 Difficulty-aware Progressive Supervised Instruction Fine-tuning
3.4.1 Fine-tuning dataset overview
To further enhance domain-specific retrieval capabilities, DMRetriever is fine-tuned on a labeled dataset, MTT (Massive Text Triplets; Algorithm 3) comprising: DM-MTT for disaster management domain and GD-MTT for general domain. DM-MTT is obtained by applying false positive filtering (§3.4.2) and hard negative mining (§3.4.3) to query–passage pairs in DM-MTP, while pairs in GD-MTP undergo hard negative mining to produce GD-MTT. Its composition is listed in Table 11.
3.4.2 Mutual-agreement False Positive Filter
Previous studies (alberti2019synthetic; xu2024bmretriever; wang2022text) show that LLM synthetic data can be noisy, and they commonly adopt a consistency-based filtering (CBF for short) approach: discarding pairs in which LLM-generated positive passage is not ranked within the top-$k$ results returned by a reference embedding model.
Given the limited capability of existing open-source IR models in disaster management, we propose a mutual-agreement-based false positive filtering method to more effectively remove low-quality pairs from DM-MTP (A-2 in Figure 2, lines 1–8 in Algorithm 3). For each query, we use the top three open-source IR models identified in DisastIR for its corresponding search intent (Table 4) to retrieve the top-$N$ passages. A query is retained only if all three models return the same top-$N$ passages; in that case, its top-1 passage is designated as the positive passage of this query.
3.4.3 Difficulty-aware Hard Negative Mining
Query–positive passage pairs further undergo hard negative mining to provide challenging negatives for enhancing DMRetriever’s performance. Given that DMRetriever spans a wide range of model parameter sizes with different learning capacities, we propose a difficulty-aware hard negative mining strategy that caters to each model’s learning ability by providing negatives at appropriate difficulty levels.
Inspired by moreira2024nv, for a query $q_{i}$ with a positive passage $p_{i}^{+}$, a passage $p_{k}$ is treated as a hard negative $p_{i,k}^{-}$ if
$\mathrm{sim}(q_{i},p_{k})<\mathrm{sim}(q_{i},p_{i}^{+})\times\alpha.$ (3)
For each query $q_{i}$, we employ the best-performing IR model corresponding to its search intent (See Table 4) to retrieve the top 200 passages from the associated passage corpus. From this pool, we select the top-$K$ passages that satisfy Equation 3, forming the hard negative set $\mathcal{H}_{i}$ (A-3 in Figure 2, Lines 10–15 in Algorithm 3).
The margin $\alpha$ serves as a tunable parameter that controls the difficulty of mined negatives, referred to as the difficulty level. Instead of applying a uniform difficulty level across all DMRetriever variants, we adjust $\alpha$ based on model size to better align with their learning capacities. This yields datasets tailored to specific difficulty levels, denoted as MTT-$\alpha$.
Model Scale QA QAdoc TW FC NLI STS Avg.
[…]
thenlper-gte-small (li2023gte) 33M 18.04 9.13 10.95 49.63 37.51 55.55 30.14
arctic-embed-m (merrick2024arctic) 109M 33.15 14.04 8.48 35.07 38.67 56.20 30.94
thenlper-gte-base (li2023gte) 109M 9.18 5.42 37.91 60.45 42.52 46.07 33.59
arctic-embed-m-v1.5 (merrick2024arctic) 109M 25.76 30.41 17.95 47.97 42.88 64.16 38.19
arctic-embed-s (merrick2024arctic) 33M 38.58 28.81 21.33 47.21 39.85 66.96 40.46
bge-small-en-v1.5 (xiao2024c) 33M 56.91 51.19 25.15 55.17 32.87 64.54 47.64
bge-base-en-v1.5 (xiao2024c) 109M 51.50 52.78 46.72 59.93 41.16 68.63 53.45
[…]
arctic-embed-m-long (merrick2024arctic) 137M 21.51 10.86 19.24 36.13 41.67 54.94 30.73
arctic-embed-l (merrick2024arctic) 335M 40.56 30.19 14.98 32.64 34.20 56.10 34.78
bge-large-en-v1.5 (xiao2024c) 335M 56.76 54.45 32.20 54.90 35.11 64.47 49.65
gte-base-en-v1.5 (li2023gte) 137M 60.51 55.62 46.26 52.24 39.59 70.40 54.10
mxbai-embed-large-v1 (emb2024mxbai) 335M 64.24 62.63 39.94 58.12 40.18 68.01 55.52
arctic-embed-m-v2.0 (merrick2024arctic) 305M 61.22 62.20 47.01 57.79 42.29 64.51 55.84
[…]
arctic-embed-l-v2.0 (merrick2024arctic) 568M 55.23 59.11 38.11 60.10 41.07 62.61 52.70
gte-large-en-v1.5 (li2023gte) 434M 67.37 58.18 39.43 52.66 34.45 66.47 53.09
Qwen3-Embedding-0.6B (qwen3embedding) 596M 66.10 52.31 62.38 64.89 50.30 67.39 60.56
mulling-e5-large-instruct (wang2024multilingual) 560M 67.97 64.64 62.25 66.78 48.51 63.42 62.26
mulling-e5-large (wang2024multilingual) 560M 66.99 64.01 62.81 59.87 50.93 74.12 63.12
gte-Qwen2-1.5B-instruct (li2023gte) 1.5B 69.85 59.17 65.09 62.73 55.51 73.58 64.32
inf-retriever-v1-1.5b (infly-ai_2025) 1.5B 69.41 64.29 62.99 65.39 54.03 73.92 65.01
[…]
Qwen3-Embedding-8B (qwen3embedding) 7.6B 44.21 34.38 41.56 42.04 32.53 42.95 39.61
gte-Qwen2-7B-instruct (li2023gte) 7.6B 70.24 47.41 63.08 31.62 53.71 74.88 56.82
NV-Embed-v1 (moreira2024nv) 7.9B 68.06 62.70 56.02 59.64 48.05 67.06 60.26
Qwen3-Embedding-4B (qwen3embedding) 4B 67.20 59.14 65.28 67.16 53.61 58.51 61.82
e5-mistral-7b-instruct (wang2023improving) 7.1B 65.57 64.97 63.31 67.86 47.55 66.48 62.58
NV-Embed-v2 (lee2024nv) 7.9B 74.47 69.37 42.40 68.32 58.20 76.07 64.80
inf-retriever-v1 (infly-ai_2025) 7.1B 72.84 66.74 66.23 65.53 51.86 75.98 66.53
SFR-Embedding-Mistral (SFRAIResearch2024) 7.1B 71.41 67.14 69.45 70.31 50.93 72.67 66.99
Linq-Embed-Mistral (LinqAIResearch2024) 7.1B 74.40 70.31 64.11 70.64 52.46 71.25 67.19
[…]
Table 2: Experiment results across all six search intents at various scales in DisastIR-Test. Bold and underline denote the best and second-best scores within each size group. indicates improvement over the best baseline is statistically significant with $p$-value <0.05<0.05 evaluated using the one-tailed Wilcoxon signed-rank test.
Size QA QAdoc TW FC NLI STS Avg
33M 57.72 55.83 57.40 59.17 41.55 64.71 56.06
109M 59.41 58.24 58.72 60.68 42.13 62.93 57.02
335M 64.10 61.36 62.09 63.33 44.86 66.55 60.38
596M 71.39 67.00 67.23 68.71 49.91 71.66 65.98
4B 72.35 68.68 70.06 71.09 53.69 74.47 68.39
7.6B 73.47 69.77 69.42 71.44 53.91 75.56 68.93
Table 3: Performance of DMRetriever-PT.
3.4.4 Progressive Supervised Instruction Fine-tuning
Previous studies show that retrieval performance degrades when trained with negatives that are either too difficult or too simple (merrick2024arctic). To address this, for small-sized variants of DMRetriever (33M, 109M, and 335M) that have limited learning capacity after pre-training, we introduce progressive instruction fine-tuning (B-3 in Figure 2, Lines 9–13 in Algorithm 1), inspired by curriculum learning (bengio2009curriculum). This approach gradually increases the difficulty of training samples across iterations, [^3] An iteration refers to a curriculum stage in Lines 9–12 of Algorithm 1, each consisting of multiple training batches. where the best validation checkpoint from one iteration initializes the next. At each stage, the difficulty level $\alpha$ is raised to make negatives harder, with models progressively trained on MTT-0.65, 0.75, and 0.85.
In contrast, the larger variants (596M, 4B, and 7.6B), which already exhibit strong performance after pre-training, are fine-tuned only on MTT-0.95. $\alpha$ value selections are detailed in Appendix D.
To further adapt DMRetriever’s embeddings to different search intents, we perform instruction fine-tuning, where each query $q_{i}$ is prepended with an intent-specific instruction $I_{q}$, forming the input $\tilde{q}_{i}=[I_{q};q_{i}]$. The model is then fine-tuned on MTT-$\alpha$ at each iteration using the InfoNCE loss:
$\mathcal{L}_{\text{ft}}=-\log\frac{e^{\mathrm{sim}(\tilde{q}_{i},p_{i}^{+})}}{e^{\mathrm{sim}(\tilde{q}_{i},p_{i}^{+})}+\sum\limits_{p_{i,k}^{-}\in\mathcal{H}_{i}}e^{\mathrm{sim}(\tilde{q}_{i},p_{i,k}^{-})}}.$ (4)
4 Experiment Setup
4.1 DisastIR-DevLite and DisastIR-Test Construction
Prior works (boteva2016full; yang2018hotpotqa; thakur2021beir) typically construct validation sets by randomly splitting queries while keeping the passage corpus identical across validation and test sets, which makes repeatedly encoding the full passage corpus computationally expensive.
To address this, we introduce DisastIR-DevLite (See Algorithm 4), a lightweight validation set derived from DisastIR [^4] DisastIR is the only publicly released IR benchmark for disaster management available at this repository. by sampling queries and constructing a much smaller passage corpus tailored to them, while the remaining queries form DisastIR-Test with the full DisastIR passage corpus. Full construction details are provided in Appendix E.
[…]
4.2 Baselines and Metrics
Baselines are drawn from two sources: (1) models evaluated in DisastIR, and (2) the latest high-performing open-source models on the MTEB retrieval benchmark  [^5] As of August 15, 2025. (muennighoff2022mteb) that have not been evaluated on DisastIR.
Some baselines in DisastIR are fine-tuned with knowledge distillation (KD), which uses external supervision from a teacher model. In such cases, models are not trained solely on ground-truth data but are also guided by teacher’s high-quality interpretation of it. We thus exclude these KD-based models as unfair comparisons, leaving 29 baselines ranging from 33M to 7.9B parameters. Detailed reasons for this exclusion are in Appendix G. Baseline implementations are detailed in Appendix H.
Model performance is evaluated using Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) on DisastIR-Test. DMRetriever training setting is detailed in Appendix I.
Figure 3: Effects of different fine-tuning datasets. DM-MTT and GD-MTT are datasets used during fine-tuning from disaster management and general domains. Effects of different pre-training datasets are shown in Figure 7.
5 Experiment Results and Analyses
We investigate three research questions: RQ1: How does DMRetriever perform across scales compared to existing baselines? RQ2: Can our data refinement pipeline generate high-quality data for model learning? RQ3: Can our three-stage training framework enable models to effectively absorb intensive domain knowledge?
5.1 DMRetriever Performance (RQ1)
Table 3.4.3 details DMRetriever performance against all baselines across all six search intents. Across all scales, our model ranks first on every individual intent and surpasses the previous SOTA by up to 7 points in average performance (335M variant). This performance culminates in our largest model, DMRetriever-7.6B, which sets a new SOTA on the benchmark with an average score of 71.37, surpassing the previous best by 4.3 points.
[…]
5.2 DMRetriever-PT Performance (RQ1)
We further evaluate DMRetriever-PT (DMRetriever-Pre-Training), which uses only unlabeled data for pre-training (§3.3.2). As shown in Table 3, it outperforms many supervised models reported in Table 3.4.3. Notably, 4B-PT surpasses all supervised baselines, while 596M-PT achieves comparable performance.
This strong unsupervised performance is especially valuable in disaster management, where labeled data is costly and scarce, making our approach a practical and data-efficient way to rapidly develop high-quality IR models for this domain.
5.3 Ablations on Data Refinement Pipeline (RQ2)
Ablations on Training Dataset Component
Figures 3 and 7 highlight the importance of disaster management-specific and general-domain corpora in shaping model performance. General-domain data introduces diverse language patterns and semantic relevance, enhancing model robustness and adaptability (xu2024simrag).
Disaster management corpus grounds model in the context of this domain, enhancing its ability to adapt to domain-specific scenarios.
This yields substantial improvements, particularly on search intents that demand intensive domain knowledge (e.g., NLI and Twitter) (yin2025disastir), with gains of up to 2.9 points on Twitter in 596M variant. These results underscore the importance of disaster management-specific data for effective domain adaptation and model performance.
Figure 4: Ablation study of different false positive filtering methods. NoPos: No false positive filtering. CBF: Consistency-based filtering. T1–3: Top-NN mutual-agreement filtering strategy, where $N=1,2,3$.
Figure 5: Ablations of progressive instruction fine-tuning. DirectIt3, BiAttn-Ada, DM-PT, and Ins mean fine-tuning directly using iteration 3 data (MTT-0.85), Bidirectional attention adaptation, DMRetriever model, and instruction, respectively. Breakdown for each search intent is in Figures 9, 10, and $1$.
Ablations on False Positive Filter Methods
We further ablate our data refinement pipeline to assess its role in improving data quality and model performance. We begin with false positive filter, comparing (1) no filtering (NoPos) and (2) a consistency-based filter (CBF; details in Appendix J). Difficulty-aware hard negative mining is analyzed together with three-stage training framework in §5.4.
As shown in Figure 4, CBF method yields only marginal gains over NoPos, improving average performance by 0.4 and 0.5 points for 33M and 596M variants. In contrast, our top-$N$ mutual-agreement filtering consistently outperforms both baselines across most search intents and in average performance, achieving up to 2.2 points higher average performance. This shows CBF struggles to filter low-quality pairs in disaster management, mainly due to the weaknesses of individual IR models in this domain, while our mutual-agreement approach addresses this limitation through model consensus.
An appropriate choice of $N$ further enhances performance: a less strict top-$1$ agreement ($N=1$) may retain noisy pairs, while a stricter top-3 agreement ($N=3$) may exclude true positives, slightly affecting performance.
5.4 Ablations on Three-stage Training Frameworks (RQ3)
Ablations on Progressive Instruction Fine-tuning
As shown in Figure 5a, progressive fine-tuning is essential for model performance. Skipping intermediate stages and directly fine-tuning DMRetriever-PT on the final stage dataset (details in Appendix K) causes a sharp performance drop, with up to 1.4 points (33M variant) lower than its pre-trained variant. This degradation indicates a mismatch between model capacity and data difficulty, as the final-stage negatives are too challenging for the model’s early-stage learning.
In contrast, our progressive method, which gradually increases the difficulty of hard negatives, enables steady improvement across iterations (Figure 8). This confirms the necessity of progressive fine-tuning and highlights the difficulty-aware hard negative mining that enables progressive training by providing controlled negatives.
We also find that adding intent-specific instructions to query significantly improves performance (Figure 5c), with gains of up to 2.3 points for 109M variant. This suggests that instruction tuning helps model better align its embeddings with diverse search intents (asai2022task; wang2023improving).
Ablations on Pre-training
Adaptation to bidirectional attention is vital for decoder-only backbones, yielding a significant performance improvement (Figure 5b) with up to 2.7 points improvement (596M variant). This indicates that bidirectional attention could yield more effective embeddings than causal attention, as it allows model to capture knowledge from future tokens (lee2024nv).
Contrastive pre-training is crucial for effective domain knowledge adaptation, especially for smaller models, yielding robust initialization for retrieval tasks (Figure 5c). The performance gains, however, decrease with increasing model size, as larger models may already possess the capacity to capture domain knowledge (xu2024bmretriever).
6 Conclusion
We present DMRetriever, a family of six dense retrieval models (33M to 7.6B) designed to advance information access in disaster management. DMRetriever is trained through a novel three-stage training framework with training data produced by an advanced data refinement method.
[…]
Limitations
While DMRetriever achieves SOTA performance across all search intents and scales, several aspects merit further investigation. The current work focuses primarily on textual retrieval; extending it to multi-modal retrieval would enhance its applicability in real-world disaster scenarios. In addition, DMRetriever is focused on English-language resources, and future work could explore multilingual and cross-lingual retrieval to broaden its global utility.
Ethics Statement
DMRetriever is designed to support disaster management by enhancing retrieval performance across diverse search intents. All training data are drawn from publicly available sources, with no personally identifiable information included. We acknowledge potential risks of misuse, such as the amplification of rumors during disasters. To mitigate such concerns, DMRetriever is released strictly for research purposes.
A potential issue is test set contamination, where some test samples may overlap with training data (sainz2023nlp). We conduct detailed data contamination analysis in Appendix C, where we do not observe any overlap between train and test sets.
Acknowledgments
This work used DeltaAI and Delta GPU at the National Center for Supercomputing Applications through allocation CIV250019, CIS250751, and CIV250021 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Input:
Backbone $\mathcal{M}$ (encoder- or decoder-only); MLM training data $\mathcal{D}_{MLM}$; Difficulty level schedule $\{\alpha_{t}\}$ for progressive fine-tuning
Output: Final DMRetriever model
1 Stage 1: Bidirectional attention adaptation (decoder-only backbone)
2 if $\mathcal{M}$ is decoder-only then
3   Replace causal mask with full attention mask;
4   Train $\mathcal{M}$ on $\mathcal{D}_{MLM}$ using the MLM objective
5Stage 2: Unsupervised contrastive pre-training
6 Construct $\mathrm{MTP}$ via Algorithm 2;
7 Pre-train $\mathcal{M}$ on $\mathrm{MTP}$ using InfoNCE loss (Eq. 2) with in-batch negatives
8 Stage 3: Difficulty-aware Progressive Supervised Instruction Fine-tuning
9 foreach iteration $t$ do
10   Construct $\mathrm{MTT}\text{-}\alpha_{t}$ via Algorithm 3 with difficulty level $\alpha_{t}$;
11   Fine-tune $\mathcal{M}$ on $\mathrm{MTT}\text{-}\alpha_{t}$ using supervised InfoNCE loss (Eq. 4);
12   Update $\mathcal{M}$ with the best validation checkpoint
return $\mathcal{M}$
Algorithm 1 Three-stage training pipeline of DMRetriever
Input:
$\mathcal{C}_{GD}$: general-domain corpora; $\mathcal{R}_{dm}$: raw disaster-management documents (e.g., PDFs); $\mathcal{I}$: search intents; $\Gamma_{i}$: heuristic rule for intent $i$; $\mathcal{E}^{need}_{i},\mathcal{E}^{query}_{i}$: in-context examples for information-need generation, and for query–positive passage generation;
$\mathrm{Prompt}^{need},\mathrm{Prompt}^{query}$: prompt templates for information-need generation, and for query–positive passage generation
Output:
$\mathrm{MTP}=\mathrm{GD\mbox{-}MTP}\cup\mathrm{DM\mbox{-}MTP}$
1
21. Build DM-MTP:
3 foreach $d\in\mathcal{R}_{dm}$ do
4 $\mathcal{C}_{dm}\leftarrow\mathrm{ProcessToPassages}(d)$;
5
6foreach $p\in\mathcal{C}_{dm}$ do
7 $s\leftarrow\mathrm{LLM}(p,\mathcal{I},\mathcal{E}^{need}_{i},\mathrm{Prompt}^{need})$;
8 $(q,p^{+})\leftarrow\mathrm{LLM}(s,p,\mathcal{E}^{query}_{i},\mathrm{Prompt}^{query})$;
9 $\mathrm{DM\mbox{-}MTP}\leftarrow\mathrm{DM\mbox{-}MTP}\cup\{(q,p^{+})\}$;
10
11
122. Build GD-MTP:
13 foreach $r\in\mathcal{C}_{GD}$ do
14   Identify intent $i$ (by dataset design);
15 $(q,p^{+})\leftarrow\mathrm{ApplyHeuristic}(r,\Gamma_{i})$;
16 $\mathrm{GD\mbox{-}MTP}\leftarrow\mathrm{GD\mbox{-}MTP}\cup\{(q,p^{+})\}$;
17
18
193. Combine and output:
$\mathrm{MTP}\leftarrow\mathrm{GD\mbox{-}MTP}\cup\mathrm{DM\mbox{-}MTP}$
Algorithm 2 Construction of MTP
QA QAdoc Twitter FactCheck NLI STS
Top 1 NV-Embed-v2 Linq-Emb-Mis SFR-Emb-Mis Linq-Emb-Mis NV-Embed-v2 NV-Embed-v2
Top 2 Linq-Emb-Mis NV-Embed-v2 inf-retriever-v1 SFR-Emb-Mis gte-Qwen2-1.5B-ins inf-retriever-v1
Top 3 inf-retriever-v1 SFR-Emb-Mis Linq-Emb-Mis NV-Embed-v2 inf-retriever-v1-1.5b gte-Qwen2-7B-ins
Table 4: Top-3 IR models across different search intents identified in DisastIR. Linq-Emb-Mis and SFR-Emb-Mis are shorten for Linq-Embed-Mistral and SFR-Embedding-Mistral
Task Instruction Format
QA Given the question, retrieve most relevant passage that best answers the question
QAdoc Given the question, retrieve most relevant document that answers the question
Twitter Given the user query, retrieve the most relevant Twitter text that meets the request
FactCheck Given the claim, retrieve most relevant document that supports or refutes the claim
NLI Given the premise, retrieve most relevant hypothesis that is entailed by the premise
STS Given the sentence, retrieve the sentence with the same meaning
Table 5: Instruction formats used for each search intent.
Input:
$\mathcal{Q}_{GD}$: queries from GD-MTP (with $(q,p^{+})$ pairs); $\mathcal{Q}_{DM}$: queries from DM-MTP; NN: top-NN passages for mutual-agreement; $\mathcal{M}_{top3}$, $\mathcal{M}_{best}$: top-3 and best IR models per intent (from DisastIR); $\alpha$: difficulty level for hard negative selection
Output: $\mathrm{MTT}\text{-}\alpha=\mathrm{GD\mbox{-}MTT}\text{-}\alpha\cup\mathrm{DM\mbox{-}MTT}\text{-}\alpha$
1
21. Mutual-agreement false positive filter
3 foreach $q\in\mathcal{Q}_{DM}$ do
4   Identify intent $i$;
5 foreach $m\in\mathcal{M}_{top3}(i)$ do
6      Retrieve top-NN passages $R_{m}(q)$;
7
8 $\mathcal{C}_{q}\leftarrow\bigcap_{m\in\mathcal{M}_{top3}(i)}R_{m}(q)$;
9 if $\mathcal{C}_{q}\neq\emptyset$ then
10      Add $(q,\text{top-1}(\mathcal{C}_{q}))$ to $\mathcal{D}^{+}_{DM}$;
11
12
[…]
14
152. Difficulty-aware hard negative mining
16 foreach $(q,p^{+})\in\mathcal{D}^{+}$ do
17   Identify intent $i$;
18   Retrieve top-200 $R_{best}(q)$ using $\mathcal{M}_{best}(i)$;
19 $\mathcal{H}_{q}\leftarrow\{p^{-}\in R_{best}(q)\setminus\{p^{+}\}\mid\mathrm{sim}(q,p^{-})<\mathrm{sim}(q,p^{+})\times\alpha\}$;
20   Add $(q,p^{+},\text{top-}k(\mathcal{H}_{q}))$ to $\mathrm{DM\mbox{-}MTT}\text{-}\alpha$ if $q\in\mathcal{Q}_{DM}$ else $\mathrm{GD\mbox{-}MTT}\text{-}\alpha$;
21
22
233. Combine and output: $\mathrm{MTT}\text{-}\alpha\leftarrow\mathrm{GD\mbox{-}MTT}\text{-}\alpha\cup\mathrm{DM\mbox{-}MTT}\text{-}\alpha$;
return $\mathrm{MTT}\text{-}\alpha$
Algorithm 3 Construction of MTT-$\alpha$
Appendix A Details of DM-MTP Construction
We propose a three-stage pipeline to construct DM-MTP: (1) web crawling and processing of disaster-management domain PDF files, (2) information needs generation, and (3) query–positive passage pairs generation.
For web crawling, we construct queries in the format “event type name + location name + .pdf”. We specifically target PDFs because they are more likely to contain structured, information-rich, and credible content, often originating from peer-reviewed publications or official institutions. To ensure coverage, we consider 301 disaster event types as defined in undrr2020hazard and combine them with the names of the 50 U.S. states, given both the availability of high-quality resources in the U.S. context and the fact that our model is designed for English-language data.
All collected PDF files are then processed into semantically coherent text chunks through a multi-step pipeline proposed by yin2025disastir: (1) exact-URL deduplication, where duplicate documents are removed by identifying identical download links; (2) text extraction and preprocessing, where each PDF is converted into plain text and non-textual elements such as tables and figures are removed; (3) locality-sensitive hashing (LSH) deduplication, which eliminates near-duplicate documents with overlapping content; (4) semantic chunking, where cleaned documents are segmented into passages of fewer than 256 tokens to balance retrievability and semantic integrity; and (5) embedding-based near deduplication, where dense embeddings are computed for all chunks, an ANN index is built, and pairs with cosine similarity above 0.9 are discarded. Chunks obtained are regarded as disaster management passages in DM-MTP.
In the second and third stages, given a passage and the definition of a search intent, an LLM generates information-need statement. The information statement together with the passage is used to construct a user query and its directly relevant positive passage. Prompts for query generation and relevant passage generation based on disaster management-related passages under different search intents are adapted from yin2025disastir.
In total, DM-MTP contains 3.3 million query–passage pairs, with an estimated generation cost of about \$2,000 using the GPT-4o-mini API.
Appendix B Details of GD-MTP Construction
To construct DG-MTP, we generate query–positive passage pairs from publicly available unlabeled corpora using intent-specific strategies as follows: (1) For QA, QAdoc, and TW intents, the user question serves as the query, and its corresponding answer is the positive passage. (2) For FC intent, the claim is used as the query, and a document that supports or refutes it is selected as the positive passage. (3) For NLI intent, the premise is treated as the query and the entailed hypothesis as the positive passage. (4) For STS intent, one sentence from a similar pair is randomly chosen as the query and the other as the positive passage.
Appendix C Data Contamination Analysis
Following wang2023improving, we conduct test set data contamination analysis from two aspects: query overlap and passage corpus overlap.
Query overlap.
Following wang2023improving, we perform a string-matching analysis between queries in the DisastIR-Test and those in the training sets (MTP and MTT), and find no overlap between training and test queries.
Passage corpus overlap.
While sharing the same passage corpus (e.g., DBPedia, NQ, and TriviaQA all using Wikipedia) is a standard evaluation practice in information retrieval and is not regarded as contamination (wang2023improving), we adopt a stricter criterion by performing a string-matching analysis between passages in DisastIR-Test and those in MTP and MTT, finding no overlap between the training and test corpora.
Appendix D Determination of $\alpha$ Value during Progressive Fine-tuning
We follow the ablation study results of moreira2024nv and set $\alpha=0.95$ for DMRetriever-596M, 4B, and 7.6B.
For DMRetriever-33M, 109M, and 335M, we apply progressive fine-tuning across three iterations, training each model for one epoch per iteration. We consider three candidate ranges of $\alpha$: $\{0.55,0.65,0.75\}$, $\{0.65,0.75,0.85\}$, and $\{0.75,0.85,0.95\}$. Model performance is assessed on the validation set (DisastIR-DevLite), and the range yielding the best validation performance is selected. Results show that models progressively trained on MTT-0.65, 0.75, and 0.85 could lead to the best validation performance and $\{0.65,0.75,0.85\}$ is thus selected.
Input:
DisastIR corpus $\mathcal{C}$ with labeled queries $\mathcal{Q}$ and relevance $\mathcal{Y}$; Search intents $\mathcal{T}$; Top-3 IR models $\mathcal{M}_{top3}$ per intent; $N=80$ queries per intent for DevLite, $k=10$ passages per model
Output:
DisastIR-DevLite $(\mathcal{Q}_{\text{dev}},\mathcal{C}_{\text{dev}},\mathcal{Y}_{\text{dev}})$;
DisastIR-Test $(\mathcal{Q}_{\text{test}},\mathcal{C},\mathcal{Y}_{\text{test}})$
1
2Step 1: Sample queries for DevLite and Test
3 foreach intent $t\in\mathcal{T}$ do
4   Randomly sample NN queries $\mathcal{Q}_{\text{dev}}^{(t)}\subset\mathcal{Q}^{(t)}$;
5   Let $\mathcal{Q}_{\text{test}}^{(t)}\leftarrow\mathcal{Q}^{(t)}\setminus\mathcal{Q}_{\text{dev}}^{(t)}$
[…]
7[2pt]
8Step 2: Build passage pool per DevLite query
9 foreach $q\in\mathcal{Q}_{\text{dev}}$ with intent $t$ do
10 foreach $m\in\mathcal{M}_{top3}(t)$ do
11      Retrieve $\text{TopK}_{m}(q)$ from $\mathcal{C}$
12 $\text{Pool}(q)\leftarrow\bigcup_{m\in\mathcal{M}_{top3}(t)}\text{TopK}_{m}(q)$
[…]
14[2pt]
15Step 3: Inherit relevance labels from DisastIR
16 $\mathcal{Y}_{\text{dev}}\leftarrow\mathcal{Y}|_{\mathcal{Q}_{\text{dev}}\times\mathcal{C}_{\text{dev}}}$;  $\mathcal{Y}_{\text{test}}\leftarrow\mathcal{Y}|_{\mathcal{Q}_{\text{test}}\times\mathcal{C}}$
17
return DisastIR-DevLite and DisastIR-Test
Algorithm 4 Construction of DisastIR-DevLite and DisastIR-Test
Appendix E Details of DisastIR-DevLite and DisastIR-Test Construction
To facilitate efficient yet effective model development, we construct a lightweight validation set, referred to as DisastIR-DevLite (See Algorithm 4). For each search intent, we randomly sample 80 user queries from DisastIR to form DisastIR-DevLite, while the remaining queries constitute DisastIR-Test. For each query in DisastIR-DevLite, we apply the top three IR models corresponding to its search intent, retrieving the top-10 passages from each model. The union of these top-10 passages forms a passage pool for that query. The final DisastIR-DevLite passage corpus is constructed as the union of the passage pools of all queries in DisastIR-DevLite, whereas DisastIR-Test passage corpus is the entire DisastIR passage corpus. Query–passage relevance labels for both DisastIR-DevLite and DisastIR-Test are directly inherited from the original DisastIR.
Appendix F Efficiency and Effectiveness of DisastIR-DevLite
To evaluate the efficiency and effectiveness of DisastIR-DevLite, we compare it against DisastIR-DevFull, a counterpart that shares the same queries as DisastIR-DevLite but uses the complete DisastIR corpus. For our analysis, we sampled 100 checkpoints from each model, DMRetriever-33M and DMRetriever-596M, during training, and evaluated each checkpoint on both development sets. The evaluation metrics include average embedding time, as well as average and intent-level NDCG@10 score.
[…]
Dataset #Queries #Passages
DisastIR-DevLite 480 9,316
DisastIR-Test 9,120 239,704
Table 6: Statistics of DisastIR-DevLite and DisastIR-Test.
Figure 6: Efficiency comparison between DisastIR-DevLite and DisastIR-DevFull.
Appendix G Detailed Reasons for Excluding Knowledge Distilled (KD) Baselines
For baselines (wang2022text; awasthy2025granite) that use knowledge distillation (KD) during fine-tuning, a cross-encoder is typically employed as the teacher model. This process introduces substantial external knowledge, as the teacher cross-encoder has a more powerful architecture. Unlike bi-encoders (our primary baselines and DMRetriever), cross-encoders process queries and documents jointly, allowing deep token-level interactions that produce highly accurate relevance scores. These fine-grained scores are then used as soft labels for the student model, offering a much richer and more informative training signal than the sparse binary hard labels available to other models.
Consequently, student model does not learn from ground-truth data alone, but is guided by teacher’s high-quality interpretation of that data.
For this reason, we exclude all KD-based baselines, which include E5-small, base, large-v2 (wang2022text), and granite-embedding-125m-english (awasthy2025granite), to ensure fair comparisons.
Appendix H Information on Baseline Models and Their Implementation
We adopt baseline models evaluated in DisastIR and additionally include recent models from MTEB that were not evaluated in DisastIR. They are categorized as four scales based on model parameter sizes: small (<=109M), medium (137M - 335M), large (434M- 1.5B), and extra large (XL) (>= 4B).
Detailed specifications of the selected models are provided in Table 12, while their HuggingFace links and licenses are listed in Table 13.
For each baseline model, we adhere to the official implementation guidelines to generate normalized query and passage embeddings. All evaluations are conducted in a zero-shot setting, with input sequences truncated to 512 tokens and a task-specific instruction prepended to each query.
Appendix I Implementation Details of DMRetriever Training
Backbones of DMRetriever are shown in Table 1. For bidirectional attention adaptation of decoder-only backbones, we train for 2 epochs with a global batch size of 256. The learning rates are set according to model size (see Table 7). All experiments are conducted on a single H200 GPU.
For unsupervised contrastive pre-training, each model is trained for 1 epoch. We use 3 H200 GPUs for encoder-only backbones and 4 H200 GPUs for decoder-only backbones. To optimize GPU memory consumption, we adopt DeepSpeed ZeRO Stage 1 (rasley2020deepspeed) and gradient checkpointing. Gradient accumulation is set to 1, and the per-GPU batch size is maximized to increase the diversity of in-batch negatives (bge_embedding), leading to the global batch sizes reported in Table 7.
For progressive supervised fine-tuning, we train encoder-only backbones for 1 epoch per iteration and decoder-only backbones for 1 epoch. Fine-tuning is conducted on a single H200 GPU, with gradient checkpointing disabled to improve training speed and gradient accumulation enabled. The learning rates and global batch sizes for each model variant are listed in Table 7.
Across all experiments, we use brain floating point (bfloat16) quantization and truncate input sequences to 512 tokens. The temperature value is set as 0.01. For each query, we set one positive passage and nine passages during pre-training and fine-tuning. For decoder-only backbones, we apply LoRA with rank $r=16$ and $\alpha=32$, following prior work (lee2024nv; wang2023improving; xu2024bmretriever), leaving hyperparameter tuning for future exploration.
33M 109M 335M 596M 4B 7.6B
Bidirectional Adaptation Stage
Learning Rate - - - 1e-4 5e-5 2e-5
Batch Size - - - 256 256 256
Contrastive Pre-training
Learning Rate 4e-4 3e-4 1e-4 9e-5 4e-5 2e-5
Batch Size 2400 732 384 384 192 140
Supervised Fine-tuning
Learning Rate 4e-5 1e-5 9e-6 6e-5 2e-5 1e-5
Batch Size 1024 512 512 512 128 64
Table 7: Training hyperparameters for different model sizes during the three-stage training framework.
Appendix J Implementation of Consistency-based filtering (CBF) Method
To implement the CBF method, we follow the methodology of wang2022text. Specifically, for each search intent, we use its best-performing retrieval model identified from the DisastIR benchmark as a reference (Table 4 for best-performing retrieval models). A synthetic pair is then discarded if its LLM-generated passage is not ranked within the top-2 results retrieved by this reference model.
Appendix K Implementation of Ablations of Progressive Fine-tuning
In the ablation study of progressive fine-tuning, we fine-tune DMRetriever-33M-PT, 109M-PT, and 335M-PT on MTT-0.85 for 3 epochs. All other hyperparameters, including learning rate and batch size, are kept the same as in progressive fine-tuning which are detailed in Appendix I.
Appendix L Performance of DMRetriever-KD
To establish a fair comparison with baselines like the E5 series that employ knowledge distillation (KD) during model fine-tuning and boost performance of our smaller models, we apply KD to our fine-tuned DMRetriever-33M, 109M, and 335M models. This process yields DMRetriever-KD series (see Appendix M for KD settings). As shown in Table M, our DMRetriever-KD achieves the best average performance within their respective scales, securing the top rank across the majority of search intents.
[…]
Appendix M Knowledge Distillation Settings
We apply knowledge distillation (KD) to our DMRetriever-33M, 109M, and 335M models, creating their distilled counterparts, denoted as the DMRetriever-KD series. The distillation is performed using a specially constructed dataset, MTT-0.95-KD.
This dataset is an extension of MTT-0.95, the construction of which is detailed in Section 3.4 and Algorithm 3 (Lines 10–16). To create KD version, we use the best-performing IR model, as identified in DisastIR, for each search intent as a teacher to generate query-passage similarity scores. These scores are then added to the dataset as soft labels.
The training objective for knowledge distillation, $\mathcal{L}_{\text{KD}}$, is formulated as a weighted sum of two components. It combines the contrastive loss, $\mathcal{L}_{\text{CL}}$ (from Eq. 4), on the original hard labels with a Kullback-Leibler (KL) divergence term, $\mathcal{L}_{\text{KL}}$, for distilling the soft labels from the teacher. The balance between these two is controlled by a hyperparameter $\lambda$, as shown below: L_KD = L_CL + λL_KL In this work, we set the weighting factor $\lambda$ to 1 for all distillation experiments.
Model Scale QA QAdoc TW FC NLI STS Avg
[…]
e5-base-V2 (wang2022text) 109M 65.52 62.68 57.53 61.86 45.40 74.38 61.23
e5-small-V2 (wang2022text) 33M 65.79 62.64 60.12 61.80 46.99 74.54 61.98
[…]
granite-embedding-125m (awasthy2025granite) 125M 64.58 60.73 46.47 62.45 48.05 71.61 58.98
e5-large-V2 (wang2022text) 335M 59.92 63.07 55.37 61.85 50.69 74.69 60.93
[…]
Table 8: Performance of knowledge distillation enhanced retrieval models across multiple search intents.
Model QA QAdoc TW FC NLI STS Overall
DMRetriever-33M 0.8467 0.7504 0.7326 0.8859 0.8610 0.7255 0.9008
DMRetriever-596M 0.8995 0.7937 0.8042 0.9101 0.8730 0.7725 0.9033
Table 9: Effectiveness of DisastIR-DevLite. Kendall’s $\tau$ value between DisastIR-DevLite and DisastIR-DevFull.
Figure 7: Effects of different pre-training datasets. DM-MTP and GD-MTP are datasets used during pre-training from disaster management and general domains.
Dataset Size Task Link
GD-MTT:
MSMARCO 141k QA https://huggingface.co/datasets/microsoft/ms_marco
MSMARCO-Document-Ranking 57k QAdoc https://microsoft.github.io/msmarco/Datasets
Fever 81k FactCheck https://huggingface.co/datasets/mteb/fever
Quora-duplicates 4k STS https://huggingface.co/datasets/sentence-transformers/quora-duplicates
XNLI 96k NLI https://huggingface.co/datasets/mteb/xnli
All-NLI 145k NLI https://huggingface.co/datasets/sentence-transformers/all-nli
Signal1m-generated-queries 200k* Twitter https://huggingface.co/datasets/BeIR/signal1m-generated-queries
DM-MTT 412k
Total 1.14M
Table 11: Composition of MTT used for supervised instruction fine-tuning in Stage 3. *: After applying false positive filtering, we randomly sampled 200K examples from the remaining dataset. A portion of subsets in GD-MTP undergo hard negative mining (§3.4.3) to get GD-MTT.
Model Param Size Base Embed. Arch.
Name Size Bin Model Size
Qwen3-Embedding-8B 7.6B XL Qwen3-8B 4096 decoder
Qwen3-Embedding-4B 4B XL Qwen3-4B 2560 decoder
inf-retriever-v1 7B XL gte-Qwen2-7B-instruct 3584 decoder
NV-Embed-v2 7B XL Mistral-7B-v0.1 4096 decoder
Qwen3-Embedding-0.6B 0.6B Large Qwen3-0.6B 1024 decoder
inf-retriever-v1-1.5b 1.5B Large gte-Qwen2-1.5B-instruct 1536 decoder
Linq-Embed-Mistral 7B XL E5-mistral-7b-instruct 4096 decoder
NV-Embed-v1 7B XL Mistral-7B-v0.1 4096 decoder
SFR-Embedding-Mistral 7B XL E5-mistral-7b-instruct 4096 decoder
snowflake-arctic-embed-l 335M Medium e5-large-unsupervised 1024 encoder
snowflake-arctic-embed-l-v2.0 568M Large gte-multilingual-mlm-base 1024 encoder
snowflake-arctic-embed-m-v2.0 305M Medium bge-m3-retromae 768 encoder
gte-Qwen2-7B-instruct 7B XL Qwen2-7B 3584 decoder
snowflake-arctic-embed-m-v1.5 109M Small BERT-base-uncased 768 encoder
e5-mistral-7b-instruct 7B XL Mistral-7b 4096 decoder
snowflake-arctic-embed-m 109M Small e5-unsupervised-base 764 encoder
snowflake-arctic-embed-m-long 137M Medium nomic-embed-text-v1-uns 768 encoder
bge-large-en-v1.5 335M Medium 1024 encoder
mxbai-embed-large-v1 335M Medium 1024 encoder
snowflake-arctic-embed-s 33M Small e5-unsupervised-small 384 encoder
bge-base-en-v1.5 109M Small 768 encoder
bge-small-en-v1.5 33M Small 384 encoder
multilingual-e5-large-instruct 560M Large xlm-roberta-large 1024 encoder
thenlper-gte-base 109M Small EBRT-base 768 encoder
multilingual-e5-large 560M Large xlm-roberta-large 1024 encoder
thenlper-gte-small 33M Small MiniLM-L12-H384 384 encoder
gte-Qwen2-1.5B-instruct 1.5B Large Qwen2-1.5B 1536 decoder
gte-base-en-v1.5 137M Medium EBRT-base 768 encoder
gte-large-en-v1.5 434M Large EBRT-large 1024 encoder
Table 12: Information of all baseline models. “–” means no publicly available information is available.
Model Name Link License
Qwen3-Embedding-8B https://huggingface.co/Qwen/Qwen3-Embedding-8B apache-2.0
Qwen3-Embedding-4B https://huggingface.co/Qwen/Qwen3-Embedding-4B apache-2.0
inf-retriever-v1 https://huggingface.co/infly/inf-retriever-v1 apache-2.0
NV-Embed-v2 https://huggingface.co/nvidia/NV-Embed-v2 $c$-by-nc-4.0
Qwen3-Embedding-4B https://huggingface.co/Qwen/Qwen3-Embedding-0.6B apache-2.0
inf-retriever-v1-1.5b https://huggingface.co/infly/inf-retriever-v1-1.5b apache-2.0
Linq-Embed-Mistral https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral $c$-by-nc-4.0
NV-Embed-v1 https://huggingface.co/nvidia/NV-Embed-v1 $c$-by-nc-4.0
SFR-Embedding-Mistral https://huggingface.co/Salesforce/SFR-Embedding-Mistral $c$-by-nc-4.0
snowflake-arctic-embed-l https://huggingface.co/Snowflake/snowflake-arctic-embed-l apache-2.0
snowflake-arctic-embed-l-v2.0 https://huggingface.co/Snowflake/snowflake-arctic-embed-l-v2.0 apache-2.0
snowflake-arctic-embed-m-v2.0 https://huggingface.co/Snowflake/snowflake-arctic-embed-m-v2.0 apache-2.0
gte-Qwen2-7B-instruct https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct apache-2.0
snowflake-arctic-embed-m-v1.5 https://huggingface.co/Snowflake/snowflake-arctic-embed-m-v1.5 apache-2.0
e5-mistral-7b-instruct https://huggingface.co/intfloat/e5-mistral-7b-instruct mit
snowflake-arctic-embed-m https://huggingface.co/Snowflake/snowflake-arctic-embed-m apache-2.0
snowflake-arctic-embed-m https://huggingface.co/Snowflake/snowflake-arctic-embed-m-long apache-2.0
granite-embedding-125m-english https://huggingface.co/ibm-granite/granite-embedding-125m-english mit
bge-large-en-v1.5 https://huggingface.co/BAAI/bge-large-en-v1.5 apache-2.0
snowflake-arctic-embed-s https://huggingface.co/Snowflake/snowflake-arctic-embed-s mit
bge-base-en-v1.5 https://huggingface.co/BAAI/bge-base-en-v1.5 mit
bge-small-en-v1.5 https://huggingface.co/BAAI/bge-small-en-v1.5 mit
multilingual-e5-large-instruct https://huggingface.co/intfloat/multilingual-e5-large-instruct mit
thenlper-gte-base https://huggingface.co/thenlper/gte-base mit
multilingual-e5-large https://huggingface.co/intfloat/multilingual-e5-large-instructt apache-2.0
thenlper-gte-small https://huggingface.co/thenlper/gte-small mit
gte-Qwen2-1.5B-instruct https://huggingface.co/Alibaba-NLP/gte-Qwen2-1.5B-instruct mit
gte-base-en-v1.5 https://huggingface.co/Alibaba-NLP/gte-base-en-v1.5 apache-2.0
gte-large-en-v1.5 https://huggingface.co/Alibaba-NLP/gte-large-en-v1.5 mit
Table 13: HuggingFace model links and licenses for all evaluated models.
Figure 8: Model performance during different iterations of progressively supervised instruction fine-tuning
Figure 9: Comparisons of model performance under progressive fine-tuning and direct iteration-3 fine-tuning across 6 search intents.
Figure 10: Model performance comparisons between with and without bidirectional attention adaptation across 6 search intents.
Figure 11: Model performance under ablations of whether conducting pre-training and adding instruction across 6 search intents.